OpenStudy (kmullis6):

PLEASE HELP! /: Evaluate the following expression using the values given: Find 2r^2 + 3s^3 - r^2 + 4t^2 - r^2 if r = -2, s = -3, and t = 5. a) -19 b) -37 c) 27 d) 19

OpenStudy (kmullis6):

I came out with the answer 127 and there isn't any other answers.. I'm so confused. I suck at math, so I wouldn't be surprised if I did the problem wrong.

OpenStudy (the_fizicx99):

Plug in those values into r, s and t \(\ \color{greenyellow}{R = -2} \) \(\ \color{#19FF00}{S = -3} \) \(\ \color{aqua}{ T = 5} \) \(\ 2\color{yellowgreen}{r}\)^2 + \(\ 3\color{#19FF00}{s}^3 \) - \(\ \color{yellowgreen}{r}^2 \)^2 + 4\(\ \color{aqua}{t}^2\) - \(\ \color{yellowgreen}{r}^2\)

OpenStudy (anonymous):

uuummm.... explain ^^^^

OpenStudy (kmullis6):

@tHe_FiZiCx99 this is what I did... 2(-2)^2+3(-3)^2-(-2)^2+4(5)^2-(-2)^2 2(4)+3(9)-4+4(25)-4 8+27-4+100-4 127

OpenStudy (anonymous):

its suppoed so be 3s^3 not squared

OpenStudy (anonymous):

so r = -2 , s = -3 , t = 5 2*(-2)^2 + 3*(-3)^3 - (-2)^2 + 4*(5)^2 - (-2)^2 = 8 - 81 - 4 + 100 - 4 = 108 - 89 = 19

OpenStudy (anonymous):

a more simpler way would be to simplify the original equation first \[2r^2 + 3s^3 - r^2 + 4t^2 - r^2\] \[2r^2 + 3s^3 - 2r^2 + 4t^2\] \[3s^3 + 4t^2\] and plug in s and 3 to get your answer

OpenStudy (anonymous):

t sorry

OpenStudy (anonymous):

hey @jayz657 why does your answer always show [Math Processing Error]?

OpenStudy (bibby):

Switch browsers

OpenStudy (anonymous):

it looks fine for me

OpenStudy (kmullis6):

@DragonSage Sorry for my dumbness, but on the question, shouldn't it be ''2*(-2)^2''... I thought ^2 meant to for instance if the question were 2^2, the answer would be 4. If that's the case in this, shouldn't the answer be 16? Because 2*(-2) = 4. And 4 times 4 is 16. Sorry if this is confusing. I'm really just trying to understand. ):

OpenStudy (anonymous):

@kmullis6 look at what i wrote and plug in the values for s and t

OpenStudy (kmullis6):

@jayz657 Did that make any sense tho or am I just stupid? lol :P

OpenStudy (anonymous):

no your right but the mistake you made earlier wasnt there it was at 3s^3 you wrote 3s^2 instead of 3s^3

OpenStudy (kmullis6):

oh ok @jayz657

OpenStudy (kmullis6):

So just to verify, the first answer would be 8 or 16? Still confused on what to do there... @jayz657

OpenStudy (bibby):

\[\huge 2r^2 =2(r^2) = 2*(-2)^2=2*4 =8\]

OpenStudy (kmullis6):

That's why I'm confused, because I was always taught for instance, if the problem was 5^2, the answer would be 10. Like 5 times 5. Do the rules of ''^'' change here? I feel so stupid? :P

OpenStudy (bibby):

I don't really get what you mean. if you mean something like:\[\huge 2*5^2 = 2*(5*5)= 2*25=50\]

OpenStudy (kmullis6):

I was taught in school that you didn't multiply like that when it came to things like 5^3, for instance. If you forgot and did 5 times 3 and put 15 down you were wrong. It was supposed to be worked out like 5 times 5 times 5 = 125. Sorry that it's confusing. It's just what I was taught. @bibby

OpenStudy (bibby):

That's what I was doing. Everything I did followed PEDMAS/PEMDAS. All I was doing was evaluating the exponent part first. So (-2)^2 = 4 and then I multiplied that by 2

OpenStudy (anonymous):

BEDMAS, sounds so much better even though it's parentheses not brackets.

OpenStudy (kmullis6):

Then, I guess I really don't know what I was doing lol... I was doing 2*(-2)^2 and getting 16 because 2*(-2) = 4. 4^2 = 16.

OpenStudy (bibby):

Ooh. That's why I separated the terms in parentheses. You can't do that. \[\large 2{(-2)}^2\] is the same thing as \[\large 2*[{(-2)}^2]\]

OpenStudy (anonymous):

Result is 19 btw

OpenStudy (anonymous):

Because I'm batman.

OpenStudy (the_fizicx99):

Yeah sorry, my walmart internet wouldn't reconnect ._.

OpenStudy (kmullis6):

I get confused with Algebra easily. I'm only on my 2nd assignment in Algebra 1, so I'm sorry for spending all this time on this lol. But what happens to the 2? @bibby @iambatman Lol :P @tHe_FiZiCx99 It's ok (:

OpenStudy (bibby):

lol I'll rewrite all the work, gimme a sec

OpenStudy (the_fizicx99):

Sowwie :< Go bibbeh!!

OpenStudy (kmullis6):

hahaha i saw that @bibby :P

OpenStudy (bibby):

the mods are also gonna see that :'(

OpenStudy (anonymous):

I missed it, share :3

OpenStudy (the_fizicx99):

Missed it D;

OpenStudy (bibby):

Find 2r^2 + 3s^3 - r^2 + 4t^2 - r^2 r = -2, s = -3, and t = 5. \[\huge 2r^2 + 3s^3 - r^2 + 4t^2 - r^2\]\[\huge 2[r^2] + 3[s^3] - r^2 + 4[t^2] - r^2\] \[\huge 2[(-2)^2] + 3[(-3)^3] - (-2)^2 + 4[5^2] - (-2)^2\] I was just being mean to fiz

OpenStudy (anonymous):

Right on

OpenStudy (kmullis6):

haha

OpenStudy (bibby):

\[\large 2[(-2)^2] + 3[(-3)^3] - (-2)^2 + 4[5^2] - (-2)^2\]

OpenStudy (the_fizicx99):

Fahk ;-; missed it D:

OpenStudy (bibby):

\[\large 2[4] + 3[-27] - [(-2)^2] + 4[25] - [4]\]

OpenStudy (kmullis6):

All of those parenthesis confuse me. Aahhh! -____-

OpenStudy (bibby):

those brackets are just like parentheses. They mean, "I need a value before you can do any math with me"

OpenStudy (the_fizicx99):

They're just put there to not mixed them up, think of them as instead a + b+ c = (a) + (b) + (c)

OpenStudy (kmullis6):

What's the difference between [ ] and ( )?

OpenStudy (the_fizicx99):

Nothing, they're all the same, too many will look confusing

OpenStudy (kmullis6):

Oh. Makes sense. @bibby @tHe_FiZiCx99

OpenStudy (anonymous):

It's just to differentiate from what's going on

OpenStudy (the_fizicx99):

Yeah ^

OpenStudy (kmullis6):

@iambatman ok thanks (:

OpenStudy (bibby):

cause something like \[\large 3(−3)^3−(−2)2\] gets confusing

OpenStudy (anonymous):

But when dealing with domain and range, then that's totally different :P

OpenStudy (bibby):

anyhow\[\large 2[4] + 3[-27] - [(-2)^2] + 4[25] - [4]\]\[\large 8 + -81 - 4 + 100 - 4\]

OpenStudy (the_fizicx99):

Go bibbbbby!!!!!

OpenStudy (anonymous):

And that's 19!

OpenStudy (anonymous):

Batman is always right

OpenStudy (bibby):

good work team

OpenStudy (anonymous):

said no one ever.

OpenStudy (the_fizicx99):

Awh bib ;)

OpenStudy (bibby):

my only regret is the medals aren't distributed evenly

OpenStudy (the_fizicx99):

♥♥

OpenStudy (kmullis6):

Aw yeah that sucks ): @bibby

OpenStudy (kmullis6):

♥♥♥

OpenStudy (anonymous):

well... that escalated quickly..

OpenStudy (kmullis6):

hahaha just a little lol :P @DragonSage

OpenStudy (bibby):

It actually escalated pretty reasonably. 1 heart per post

OpenStudy (bibby):

anyhow I'll stop spamming. Apologies if my brackets made things seem unclear. They're just a way of making things more legible.

OpenStudy (kmullis6):

No it made sense after explained. Thank you so much for helping me understand this stuff better. (: @bibby

OpenStudy (kmullis6):

And thank y'all too! @tHe_FiZiCx99 @iambatman @DragonSage

OpenStudy (kmullis6):

Do you mind helping me on one more? I think I know the answer. Just wanna make sure. @bibby

OpenStudy (kmullis6):

Evaluate the following expression using the given values: Find x - 3y if x = 3 and y = -2. -9 -3 3 9 I think the answer is 9.

OpenStudy (anonymous):

3-3(-2) = 3+6 = 9

OpenStudy (bibby):

In general, it's one problem to a question but meh \[\large x-3y\]\[\large 3-[3*(-2)]\]\[\large 3-[-6]\]9

OpenStudy (bibby):

thx bby

OpenStudy (kmullis6):

Yay! Thanks guys! @iambatman @bibby

OpenStudy (anonymous):

Bibby's a boss

OpenStudy (kmullis6):

@iambatman Can I get your help again? sorry for bothering you /:

OpenStudy (bibby):

You can post new questions and tag us all

OpenStudy (kmullis6):

Ok. I feek kinda bad tho cause yall spent forever on just that one question lol @bibby

OpenStudy (bibby):

Make it up to us by posting another question where we can get more medals :D

OpenStudy (kmullis6):

Simplify the following expression: 2x - 2y + 5z - 2x - y + 3z x - 3y + 8z 2x - 2y + 8z 3y + 8z -3y + 8z So, I worked this problem out, and at the end, I got 4x - 3y - 8z. The answer was for sure C because I looked it up. What I'm wondering, is what happened to the 4x? @bibby @tHe_FiZiCx99

OpenStudy (bibby):

I'll try color coding this

OpenStudy (anonymous):

2x - 2x - 2y - y + 5z + 2z -3y + 8z

OpenStudy (kmullis6):

That's what I did, because it's in my lesson. I arranged them exactly like that. What happened to the 4x tho? /:

OpenStudy (bibby):

2x-2x = 0

OpenStudy (anonymous):

To get 4x - 3y - 8z The equation would be 2x - 2y + 5z + 2x - y + 3z

OpenStudy (kmullis6):

OH MY GAAASHHH! I WROTE IT DOWN WRONG AGAIN! I'M SO ANGRY WITH MYSELF! Sorry guys. ):

OpenStudy (bibby):

Stop apologizing :) This happens. Everyone makes careless mistakes here and there. That's why you're doing homework. So that you don't flub up the test.

OpenStudy (anonymous):

yeah what he said ^^^

OpenStudy (kmullis6):

Haha I know, but I feel like an idiot. Math's my worst subject, and this is usually why I don't do so well in this subject.

OpenStudy (bibby):

If it makes you feel better, I'm also an idiot and slept throughout most of my high school classes and still passed

OpenStudy (anonymous):

^ me too